A driver of an ambulance going 70 km/h suddenly sees another car 35.0 m ahead. It 

takes the driver 0.62 s before he applies the brakes. Once he does begin to brake, he 

reduces speed at a rate of 13.0 m/s². 

1.1 Determine if the driver will hit the car or not.


1
Expert's answer
2021-05-31T15:13:10-0400

Gives

Ambulance speed=70km/h

"v=\frac{70\times1000}{3600}=\frac{700}{36}m/sec"


Car distance(s)=35m

Time take driver before apply brack ( t)=0.62sec

Acceleration=13.0m/"sec^2"

Distance (s)=speed"\times time"


"S=\frac{700}{36} \times0.62"

S=12.05m

Distance travelled the car

"v^2=u^2+2as"

"s=\frac{v^2-u^2}{2a}"

Put value

"s=\frac{0-19.44^2}{2\times(-13)}" =14.54m

Total Distance (s)=12.05+14.54 =26.59m

D+s= 35m

Driver does not hit the car